# Jinhong Li1    Qiuping Wang1    Shujie Han2    Patrick P. C. Lee1

### v2

@@ 5.2 Microbenchmarks
We evaluate the overhead of offloading the L2P table entries to ZNS SSDs (§3.1). We consider the same four write workloads as in Exp#8 and focus on the setting of $(N_{s},N_{l})=(2,2)$ with four open segments. We set the physical space capacity as 300 GiB and issue writes to 200 GiB of data (i.e., the overprovisioned space is 50%). We vary the maximum memory size allocated for the in-memory L2P table from 50 MiB to 200 MiB. Recall that each L2P table entry has a size of four bytes for each 4-KiB block (§3.1). Thus, the in-memory L2P table can store the entries from 50 GiB to 200 GiB of data. Note that when the in-memory L2P table is allocated with 200 MiB, it can keep all L2P table entries in memory (i.e., no offloading).
@@
Figure 15 shows the results. When the in-memory L2P table has less than 200 MiB of memory space, it needs to offload some entries onto the SSD. For 4-KiB and 64-KiB random writes, a 100-MiB L2P table has lower throughput than a 200-MiB L2P table by 59.2% and 16.8%, respectively. The reason for such high throughput drops is that under uniform random writes, the chance of not caching an L2P table entry is higher. Nevertheless, larger writes have lower degradations since ZapRAID groups the L2P table entries into mapping blocks and there will be less I/O overhead of reading L2P table entries from ZNS SSDs. For 64-KiB skewed writes, a 100-MiB L2P table only has slightly lower throughput than a 200-MiB L2P table by 4.0%, meaning that the offloading overhead becomes insignificant in skewed workloads. For 64-KiB sequential writes, a 100-MiB L2P table only has slightly lower throughput than a 200-MiB L2P table by 3.6%.
Figure 15: Exp#9 (Overhead of L2P table offloading).
5.3 Trace-Driven Experiments
Exp#10 (Cloud block storage workloads).
The original Alibaba Cloud traces cover 1,000 volumes [55]. We identify that 684 out of the 1,000 volumes have more than 60% of small write requests of no more than 4 KiB and more than 25% of large write requests of at least 16 KiB (§2.2). Among the 684 volumes, we choose 20 of them whose amounts of write traffic range between 200 GiB and 240 GiB. Note that the selected 20 volumes are dominated by writes: in 18 out of 20 volumes, more than 88.3% of requests are writes, while in all volumes, more than 67.0% of requests are writes. We conduct our evaluation over five runs; in each run, we evaluate the average throughput and 95th-percentile latency of all read and write operations when running each scheme over the 20 selected volumes.
Figures 16(a) and 16(b) show the throughput results. For a single open segment with 4-KiB and 16-KiB chunks (Figure 16(a)), the throughput of ZapRAID is 69.4% and 6.4% higher than that of ZoneWrite-Only, respectively, and is close to that of ZoneAppend-Only; the findings are consistent with those in Exp#1. For multiple open segments, the findings are also consistent with those in Figure 12(d) of Exp#7. For $(N_{s},N_{l})=(2,2)$, all three schemes have similar throughput, while for $(N_{s},N_{l})=(1,3)$, ZapRAID has 25.3% higher throughput than ZoneWrite-Only and similar throughput to ZoneAppend-Only.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Throughput (single segment) (b) Throughput (multiple segments) (c) 95p latency (single segment) (d) 95p latency (multiple segments)
Figure 16: Exp#10 (Cloud block storage workloads). Average throughput and 95 latency results of all 20 volumes in Alibaba Cloud traces.
Figures 16(c) and 16(d) show the 95th-percentile latencies, in which the trends are also consistent with Exp#1 and Exp#7, respectively. For a single open segment with 4-KiB and 16-KiB chunks (Figure 16(c)), ZapRAID has 56.6% lower and 39.6% higher 95th-percentile latencies than ZoneWrite-Only under the chunk sizes of 4 KiB and 16 KiB, respectively, and comparable 95th-percentile latencies with ZoneAppend-Only. For multiple open segments (Figure 16(d)), all three schemes have similar latencies for $(N_{s},N_{l})=(2,2)$, while ZoneWrite-Only has a higher 95th-percentile latency for $(N_{s},N_{l})=(1,3)$ as its performance is bottlenecked by the small writes that are forwarded to the only small-chunk open segment.
Volume No. 1 2 3 4 5 6 7 8 9 10
Large writes 1.63 1.64 1.67 1.70 1.75 1.76 1.77 1.78 1.83 1.86
Thpt (MiB/s) (($N_{s},N_{l}$)=(1,3)) ZoneWrite-Only 2,486.4 2,341.2 2,323.1 2,401.3 2,403.7 2,333.7 2,311.0 2,360.3 2,344.8 2,410.5
Volume No. 11 12 13 14 15 16 17 18 19 20
Large writes 1.92 1.95 3.42 3.70 4.49 10.33 16.83 20.95 21.11 24.89
Thpt (MiB/s) (($N_{s},N_{l}$)=(1,3)) ZoneWrite-Only 2,372.7 2,399.1 2,412.2 2,439.7 2,593.6 2,637.1 2,703.2 2,773.4 2,862.1 2,825.4
Table 2: Exp#10 (Cloud block storage workloads). Statistics of the 20 individual volumes in Alibaba Cloud traces.
We also show the statistics of individual volumes by presenting the ratios of numbers of small writes (i.e., 4 KiB) and large writes (i.e., at least 16 KiB) over the total number of write requests as well as the throughput of each volume. For brevity, since all three schemes have close throughput under one single open segment with 16-KiB chunks and four open segments with $(N_{s},N_{l})=(2,2)$, we only present the results of one single open segment with 4-KiB chunks and four open segments with $(N_{s},N_{l})=(1,3)$. Also, we only focus on the results under ZoneWrite-Only and ZapRAID to show their variations across the volumes.
Table 2 shows the results. We sort the volumes by the ratio of large write requests. For one single open segments under 4-KiB chunks, the throughput is close over all volumes for each scheme. For four open segments with $(N_{s},N_{l})=(1,3)$, ZoneWrite-Only shows lower throughput for a lower ratio of large writes; for example, the throughput results of Volumes 2 and 20 are 2,341.2 MiB/s and 2,825.4 MiB/s with large-write ratios of 1.64% and 24.89%, respectively. The reason is that with fewer large writes, more write requests are written to the only $N_{s}=1$ small-chunk open segment, and ZoneWrite-Only fails to utilize intra-zone parallelism to issue writes efficiently. In contrast, ZapRAID maintains higher throughput with an increase of 14.7-40.8% compared with ZoneWrite-Only under $(N_{s},N_{l})=(1,3)$, across all 20 volumes.
5.4 Real-Application Experiments
We finally evaluate ZapRAID when it hosts real applications and compare it with ZoneWrite-Only and ZoneAppend-Only (note that similar application benchmarks are also found in [45]). Recall that ZapRAID achieves a high write performance gain, especially in write-intensive workloads dominated by small writes (e.g., by up to 77.2% higher write throughput than ZoneWrite-Only from Exp#1), while maintaining read performance (from Exp#2). In real application workloads mixed with reads and writes, ZapRAID may not exhibit the same significant write-only performance gain as in Exp#1, yet its performance improvement remains observable.
To host Linux applications, we export our ZapRAID prototype as an NVMe-over-Fabrics (NVMe-oF) target with TCP transport [10] and deploy the NVMe-oF host in the same testbed machine, so that the prototype serves as a Linux block device. We allocate the block device with a logical space capacity of 1 TiB, format the block device as an EXT4 file system, and run different types of applications atop the file system. As in §5.3, we consider the four settings of open segments. Note that the amount of write traffic may exceed the logical space capacity. Thus, we reserve an additional 25% of the physical space capacity for garbage collection.
Exp#11 (File writes).
We evaluate ZapRAID in file writes using Sysbench (v1.0.20) [12] to generate file write patterns. We first generate 128 files of 800 MiB each (i.e., totaling 100 GiB), and then use Sysbench to issue continuous 4-KiB write requests to the files for 10 minutes. We use one thread to issue writes in asynchronous mode with direct I/O. We issue fsync at the end of each run.
Figure 18 shows the write throughput results. For a single open segment (Figure 18(a)), ZapRAID benefits from intra-zone parallelism, achieving 38.6% higher write throughput for 4-KiB chunks than ZoneWrite-Only, and similar write throughput for 16-KiB chunks. ZapRAID also matches the write throughput of ZoneAppend-Only. The write throughput of the three schemes and the performance gains of ZapRAID are not as high as in Exp#1, as the workload contains frequent small writes, which necessitate frequent file system metadata updates with high synchronization overhead and limited write concurrency. For multiple open segments (Figure 18(b)), all three schemes show similar write throughput for $(N_{s},N_{l})=(2,2)$. ZapRAID achieves 20.2% higher throughput than ZoneWrite-Only for $(N_{s},N_{l})=(1,3)$; the findings are similar to those in Exp#7 under small writes (Figure 12(a)), albeit with a lower improvement.
Exp#12 (Relational databases).
We evaluate ZapRAID when it hosts MySQL (v8.0.30) [3], which acts as a relational database, and run the TPC-C benchmark [13]. We modify some configurations of MySQL for better performance. Specifically, we set the maximum binary log size as 2 GiB and the buffer pool size as 32 GiB. We also configure MySQL to flush the transaction log once per second and use direct I/O in flushing. We first load 1,000 TPC-C warehouses to MySQL, with a total of around 100 GiB data, and then run the TPC-C benchmark for one hour with 32 connections.
Figure 18 depicts the throughput in transactions per second (TPS). For a single open segment with 4-KiB chunks (Figure 18(a)), ZapRAID shows 16.0% higher write throughput than ZoneWrite-Only and similar throughput to ZoneAppend-Only. The limited improvement over ZoneWrite-Only is due to the low number of concurrent write requests and around 55% of all requests are reads, which are not optimized by Zone Append. For other settings of open segments, all three schemes have similar performance.
[Uncaptioned image] [Uncaptioned image]
(a) Single segment (b) Multiple segments
Figure 17: Exp#11 (File writes).
[Uncaptioned image] [Uncaptioned image]
(a) Single segment (b) Multiple segments
Figure 18: Exp#12 (Relational databases).
Exp#13 (Key-value stores).
We evaluate ZapRAID when it hosts RocksDB (v9.7.2) [7] and use the benchmarking tool db_bench [5] to issue key-value pairs; RocksDB is an LSM-tree-based [63] key-value store that issues large sequential I/Os. We focus on two workloads: fillrandom (i.e., filling an empty database with randomly ordered keys) and overwrite (i.e., randomly overwriting an existing database). We set the total number of key-value pairs as 12.5 million with a value size of 8,000 bytes. We disable compression, and use eight threads to issue requests. For each run, we reset the volume, issue the fillrandom workload, and finally issue the overwrite workload.
Figures 20 and 20 show the throughput in number of operations per second (KOPS). For the fillrandom and overwrite workloads under a single open segment with 4-KiB chunks (Figures 20(a) and 20(a)), ZapRAID achieves 22.4% and 24.3% higher throughput than ZoneWrite-Only, respectively. For other settings, the throughput of all three schemes is similar. The improvement of ZapRAID is less pronounced than in microbenchmarks (Exp#1 and Exp#7), as both fillrandom and overwrite workloads have significant read traffic (about 50% of all I/O traffic) from frequent LSM-tree compaction operations, which involve retrieving and rewriting key-value pairs; Zone Append does not optimize reads.
[Uncaptioned image] [Uncaptioned image]
(a) Single segment (b) Multiple segments
Figure 19: Exp#13 (Key-value stores). Throughput of the fillrandom workload.
[Uncaptioned image] [Uncaptioned image]
(a) Single segment (b) Multiple segments
Figure 20: Exp#13 (Key-value stores). Throughput of the overwrite workload.
Exp#14 (Concurrent applications).
We combine two types of write-intensive workloads with different write patterns to evaluate ZapRAID. Specifically, we use Sysbench [12] to generate file write patterns with small writes. We also host MySQL and run the oltp_write_only benchmark, which issues large writes with less than 10% of reads. Our setting mimics a virtualized environment that hosts multiple workloads simultaneously in an SSD RAID array [30]. For small file writes, we follow the same approach as in Exp#11. For the oltp_write_only benchmark, we use Sysbench to fill the MySQL database with 8 tables and 10 million rows, and then run DELETE, INSERT, and UPDATE queries with 64 threads. Before each run, we clear and reformat the block device as an ext4 file system, and then execute both benchmarks simultaneously for 10 minutes.
Figure 22 shows the write throughput results for file small writes. The findings are similar to Exp#11. ZapRAID achieves 26.3% and 34.5% higher write throughput under a single open segment with 4-KiB chunks (Figure 22(a)) and four open segments with $(N_{s},N_{l})=(1,3)$ (Figure 22(b)), respectively. Figure 22 depicts the throughput of the oltp_write_only benchmark in transactions per second (TPS). For a single open segment with 4-KiB chunks (Figure 22(a)), ZapRAID achieves 32.9% higher write throughput than ZoneWrite-Only and has similar throughput to ZoneAppend-Only. For other settings of open segments, all three schemes show similar performance.
[Uncaptioned image] [Uncaptioned image]
(a) Single segment (b) Multiple segments
Figure 21: Exp#14 (Concurrent applications). Throughput of small file writes.
[Uncaptioned image] [Uncaptioned image]
(a) Single segment (b) Multiple segments
Figure 22: Exp#14 (Concurrent applications). Throughput of the oltp_write_only benchmark.
@@ 6 Related Work
Storage systems for ZNS SSDs.
Several studies explore new storage system designs for ZNS SSDs. Bjorling et al. [21] adapt F2FS [48] and RocksDB [7] for ZNS SSDs, and show that the adapted systems have lower I/O amplification and higher performance than conventional SSDs. Other studies focus on the performance and management aspects of ZNS SSDs. Examples include improving the performance of log-structured merge-tree stores [63] for ZNS SSDs [40, 51, 52], optimizing host-level garbage collection [28, 33, 69, 22], proposing new ZNS interfaces for efficient zone management [59, 33, 62], enabling ZNS SSDs for swap storage [20], designing new I/O scheduling for improved intra-zone parallelism [19], extending Zone Append for sub-block data appends [66], and ensuring crash consistency on F2FS backed by ZNS SSDs [49].
While the above studies focus on a single ZNS SSD, RAIZN [45] exposes a ZNS SSD array as a single ZNS interface to applications, and focuses on fault tolerance, correctness, and crash consistency. In contrast, ZapRAID focuses on exploiting Zone Append for high performance.
Interfaces for offloading address management.
While our work focuses on the Zone Append interface, some studies propose new storage interfaces that offload address management from applications to storage devices as in Zone Append. Examples include range writes [17] and Direct File System [39] for file system management, nameless writes [77] and software-enabled flash [8] for flash-based SSDs, as well as ZEA [60] for host-managed shingled magnetic recording drives. We conjecture that ZapRAID, which builds on Zone Append, can also build on the above interfaces to improve RAID performance, and we pose this issue as future work.

### v3

@@ 5.4 Microbenchmarks (RQ3)
Exp#13 (Overhead of L2P table offloading). We evaluate the overhead of offloading the L2P table entries to ZNS SSDs (§3.1). We consider the same four write workloads as in Exp#12 and focus on the setting of $(N_{s},N_{l})=(2,2)$ with four open segments. We set the physical space capacity as 300 GiB and issue writes to 200 GiB of data (i.e., the overprovisioned space is 50%). We vary the maximum memory size allocated for the in-memory L2P table from 50 MiB to 200 MiB. Recall that each L2P table entry has a size of four bytes for each 4 KiB block (§3.1). Thus, the in-memory L2P table can store the entries from 50 GiB to 200 GiB of data. Note that when the in-memory L2P table is allocated with 200 MiB, it can keep all L2P table entries in memory (i.e., no offloading).
@@
Figure 19 shows the results. When the in-memory L2P table has less than 200 MiB of memory space, it needs to offload some entries onto the SSD. For 4 KiB and 64 KiB random writes, a 100 MiB L2P table has lower throughput than a 200 MiB L2P table by 59.2% and 16.8%, respectively. The reason for such high throughput drops is that under uniform random writes, the chance of not caching an L2P table entry is higher. Nevertheless, larger writes have lower degradations since ZapRAID groups the L2P table entries into mapping blocks and there will be less I/O overhead of reading L2P table entries from ZNS SSDs. For 64 KiB skewed writes, a 100 MiB L2P table only has slightly lower throughput than a 200 MiB L2P table by 4.0%, meaning that the offloading overhead becomes insignificant in skewed workloads. For 64 KiB sequential writes, a 100 MiB L2P table only has slightly lower throughput than a 200 MiB L2P table by 3.6%.
Figure 19: Exp#13 (Overhead of L2P table offloading).
5.5 Summary of Findings
ZapRAID is designed to integrate the strengths of both Zone Append and Zone Write primitives into a RAID system instead of always outperforming either in isolation. It is particularly effective for workloads with only small writes (e.g., 4 KiB or 8 KiB) or mixed workloads with most small writes (e.g., 4 KiB or 8 KiB) and few large writes (e.g., 16 KiB). This pattern is common in practice. For example, in the Alibaba Cloud traces (Exp#4), 684 out of 1,000 volumes (68.4%) contain more than 60% of requests no larger than 4 KiB and more than 25% of requests at least 16 KiB, with writes forming the majority of I/O.
Compared to ZoneWrite-Only, ZapRAID improves the write throughput for small writes in a single open segment, with a trade-off of introducing an extra compact stripe table. However, ZapRAID limits the size of the compact stripe table with the group-based data layout and ensures that the query overhead during degraded reads is minimal. While ZoneWrite-Only shows lower p95 latency in large writes in a single open segment (e.g., Figures 9(c) and 10(c) in Exp#5), ZapRAID introduces hybrid data management to use Zone Write for large writes and multiple open segments. When multiple open segments are available, ZapRAID shows no worse p95 write latency (Figure 16 in Exp#11).
Compared to ZoneAppend-Only, ZapRAID matches the write throughput in a single open segment because ZapRAID uses the Zone Append primitive for a single open segment. However, ZoneAppend-Only shows significantly poor degraded read performance (Exp#2) and higher memory usage for index management (Exp#3) in real-application experiments. ZapRAID outperforms ZoneAppend-Only with high degraded read performance (which is comparable to ZoneWrite-Only) and limited memory usage for index management.
We discuss the implications of our findings in terms of performance, memory usage, and functionalities of ZapRAID, compared with ZoneWrite-Only and ZoneAppend-Only.
Performance. ZapRAID demonstrates superior write performance and comparable normal read performance compared with ZoneWrite-Only and ZoneAppend-Only. Also, ZapRAID achieves significantly lower degraded read latencies than ZoneAppend-Only.
  • In terms of write performance, ZapRAID shows the highest write throughput under single-segment settings, as shown in Exp#1-Exp#4, Exp#5, and Exp#8, particularly with 4 KiB chunks, in which ZapRAID improves write throughput by up to 77.2% (Exp#5). Also, ZapRAID shows the highest write throughput and lowest write tail latencies across various configurations of $(N_{s},N_{l})$ under multiple-segment settings (Exp#1-Exp#4 and Exp#11). Furthermore, ZapRAID excels under highly concurrent write workloads (Exp#1 and Exp#10).
  • In terms of read performance, ZapRAID improves degraded read performance with lightweight stripe management. It maintains comparable degraded read performance compared with ZoneWrite-Only (Exp#6 and Exp#7) and has significantly higher degraded read performance than ZoneAppend-Only (Exp#2 and Exp#7). Note that ZapRAID is designed to maintain (rather than improve) normal read performance, so its overall throughput improvements are low for read-intensive applications (Exp#1 and Exp#4).
Memory usage. ZapRAID substantially reduces memory usage via its compact stripe table design compared with ZoneAppend-Only. It further reduces memory usage by offloading the L2P table to drives.
  • For the compact stripe table, ZapRAID reduces the compact stripe table size in two aspects. First, the group-based data layout (§3.2) significantly reduces the compact stripe table size for each zone by 66.7% (Exp#3 and Exp#5). Second, the hybrid data management (§3.3) leverages Zone Write for multiple-segment settings, so as to eliminate the need for the compact stripe table in those segments (Exp#3 and Exp#11).
  • For the L2P table, ZapRAID offloads the L2P table to drives to further reduce memory usage (Exp#13), with negligible overhead for highly skewed or sequential workloads.
Functionalities. ZapRAID supports multiple RAID schemes, crash recovery, full-drive recovery, and garbage collection, and hence adapts effectively to various scenarios.
  • ZapRAID maintains the highest write performance across various RAID schemes compared with ZoneWrite-Only and ZoneAppend-Only (Exp#8).
  • ZapRAID maintains high performance in crash recovery and full-drive recovery (Exp#9).
  • ZapRAID supports garbage collection by recycling overwritten or discarded user-written blocks. Its garbage collection overhead is small for highly skewed or sequential write workloads (Exp#12).
@@ 6 Related Work
Storage systems for ZNS SSDs. Several studies explore new storage system designs for ZNS SSDs. Bjorling et al. [5] adapt F2FS [34] and RocksDB [58] for ZNS SSDs, and show that the adapted systems have lower I/O amplification and higher performance than conventional SSDs. Other studies focus on the performance and management aspects of ZNS SSDs. Examples include improving the performance of log-structured merge-tree stores [51] for ZNS SSDs [26, 37, 38], optimizing host-level garbage collection [12, 18, 60, 6], proposing new ZNS interfaces for efficient zone management [45, 18, 49], enabling ZNS SSDs for swap storage [4], designing new I/O scheduling for improved intra-zone parallelism [3], extending Zone Append for sub-block data appends [55], and ensuring crash consistency on F2FS backed by ZNS SSDs [35].
While the above studies focus on a single ZNS SSD, RAIZN [31] exposes a ZNS SSD array as a single ZNS interface to applications, and focuses on fault tolerance, correctness, and crash consistency. BIZA [73] presents a self-governing block-interface ZNS all-flash-array that hides the ZNS interface from applications, while internally exploiting zone random write area (ZRWA) and intra-/inter-zone parallelism to improve both endurance and performance. In contrast, ZapRAID focuses on exploiting Zone Append for high performance.
Interfaces for offloading address management. While our work focuses on the Zone Append interface, some studies propose new storage interfaces that offload address management from applications to storage devices as in Zone Append. Examples include range writes [1] and Direct File System [25] for file system management, nameless writes [76] and software-enabled flash [62] for flash-based SSDs, as well as ZEA [46] for host-managed shingled magnetic recording drives. We conjecture that ZapRAID, which builds on Zone Append, can also build on the above interfaces to improve RAID performance, and we pose this issue as future work.